# False signatures of non-ergodic behavior in disordered quantum many-body systems

### v2

@@ I Introduction
The paper is organized as follows. In Sec. II, we briefly describe the peculiar properties of the probability distribution of the site-resolved magnetization defined as $\braket{n}{\hat{s}_{l}^{z}}{n}\equiv s_{l}^{z}$, where $l$ is sampled over the site indices, for the Heisenberg spin1/212-1/2- 1 / 2 chain in the presence of the quasiperiodic disorder as encountered in [44], and proceed to explore the distribution of $s_{l}^{z}$ for the Heisenberg spin1/212-1/2- 1 / 2 chain with different types of onsite disorders. This leads us to a conjecture about the relationship between the distribution of the onsite disorder and the properties of the $s_{l}^{z}$ distribution. We show the validity of this conjecture by finding the leading-order approximation for the microcanonical expectation of the site-resolved magnetization in Sec. III, and conclude that the particular structure of the $s_{l}^{z}$ distribution is an artifact of the choice of the energy window, Sec. IV. Finally, in Sec. V, we show that these specific structures have direct consequences for experiments and can be detected in the time dynamics of the $s_{l}^{z}$ expectation values.
Figure 1: The probability distribution of site-resolved magnetization, $s_{l}^{z}$, for disorder-averaged GOE matrices (a) and the HSC-QP Hamiltonian at $W=0.5$ (b) at different system sizes, $L$. For both (a) and (b) panels, $100$ eigenstates are considered at the rescaled energy $\epsilon=0.5$. The number of disorder realizations is varied, being at least $7500$ for $L\leq 16$ and $60$ for $L=18$ for panel (a), while for HSC-QP case (b) it is $10^{4}$ disorder realizations for $L\leq 16$ and $2500$ for $L=18$. The insets show the standard deviations of respective distributions, decreasing exponentially with $L$ for GOE (a) and saturating for HSC-QP model.
Figure 2: The distribution of onsite disorder is depicted for various types of disorders, including random (a), cosine (b), binary (c), tertiary (d), and triangle-wave (e). Additionally, the corresponding distributions of site-resolved magnetization are shown in (f-j), considering $100$ states near $\epsilon=0.5$ averaged over at least $200$ disorder realizations for a system of size $L=18$.
@@ II Unusual distribution for the site-resolved magnetization
We begin with a brief description of the distribution of site-resolved magnetization in a Heisenberg spin-$1/2$ chain (HSC) with quasiperiodic disorder [44]. The HSC is a paradigmatic model extensively employed to study ergodicity breaking and the interplay between disorder and quantum many-body dynamics [45]. The HSC Hamiltonian with open boundary conditions in the presence of an onsite potential reads
@@
where $\hat{s}_{l}^{\alpha}$ are spin-$1/2$ operators acting on site $l$, with $\alpha\in{x,y,z}$ and $L$ is the system size. Disorder in the model arises from the onsite potential term; initially, we shall consider an example of disorder arising from a quasiperiodic potential (QP[46]. The model has a $U(1)$ symmetry leading to magnetization, $M\equiv\sum_{l=1}^{L}\hat{s}_{l}^{z}$, being conserved. We will restrict ourselves to the magnetization-zero sector in this work. The QP introduces a deterministic, incommensurate modulation in the onsite potential, given by:
@@
where $E$ represents the energy of the state under consideration, while $E_{\min}$ and $E_{\max}$ denote the minimum and maximum eigenvalues in the spectrum, respectively. This definition ensures that $\epsilon$ is normalized to the range $[0,1]$.
Figure 3: The distribution of the difference of adjacent onsite disorder, $\delta h_{l}$, is depicted for various types of disorders, including random (a), cosine (b), binary (c), tertiary (d), and triangle-wave (e). Additionally, the corresponding distributions of differences of adjacent site-resolved magnetization, $\delta s_{l}^{z}$, (f-j), considering $100$ states near $\epsilon=0.5$, averaged over at least $200$ disorder realizations for a system of size $L=18$.
The observable of interest in this work is the site-resolved magnetization, $\hat{\mathcal{O}}\equiv\hat{s}_{l}^{z}$. In analogy with a random matrix, one could expect the observable to be Gaussian distributed around its mean value and with a standard deviation that is exponentially decreasing with system size. We confirm this expectation numerically by taking the Hamiltonian $\hat{H}$ as a matrix from the Gaussian Orthogonal ensemble [56]. The dimension of $H$ is the same as the Hilbert space dimension of HSC; the diagonal matrix elements of $H$ are taken from the Gaussian distribution $\mathcal{N}(0,1)$ with zero mean and unity variance, while the off-diagonal elements of $H$ are taken from $\mathcal{N}(0,1/\sqrt{2})$. In Figure 1(a), we show the probability distribution of site-resolved magnetization considering the expectation values $s^{z}_{l}=\braket{n}{\hat{s}_{l}^{z}}{n}$ in eigenstates |$\ket{n}$ around the rescaled energy $\epsilon=0.5$. We observe a clear single peak structure with the standard deviation decreasing exponentially with system size.
@@
    Tertiary disorder: we consider a distribution of disorder with $h_{l}$ selected randomly, with equal probability, to be $h_{l}\in\{-W,0,W\}$. This model is in the MBL regime for $W_{c}\geq 1.25\sqrt{\frac{3}{2}}$ [62]. In Fig. 2 (d,i) and Fig. 3 (d,i), we show the presence of triple peaks in the $s^{z}_{l}$ distribution, associated with the tertiary distribution of $h_{l}$, while $\delta h_{l}$ has a quinary distribution and correspondingly $\delta s_{l}^{z}$ has $5$ peaks with amplitudes decreasing with the distance from $|\delta s_{l}^{z}|=0$.
@@ III Microcanonical expectation for site-resolved magnetization
To proceed more formally, following [43] we define the Hilbert-Schmidt product in operator space as $\braket{\hat{A},\hat{B}}\equiv\frac{1}{\mathcal{D}}\mathrm{Tr}[\hat{A}^{\dagger}\hat{B}]$, where ${D}$ is the Hilbert space dimension, and the norm as $||\hat{A}||=\sqrt{\braket{\hat{A},\hat{A}}}$. The Hamiltonian of an ergodic many-body system is denoted by $\hat{H}$, and $\hat{\mathcal{O}}$ represents an arbitrary few-body observable. The operator space can be decomposed into two orthogonal subspaces: one spanned by the powers of the Hamiltonian, $\hat{H}^{k}$ (with $k\geq 0$), and an orthogonal complement. Hence, the observable $\hat{\mathcal{O}}$ can be decomposed as
@@
[…]
To establish a connection between the diagonal matrix element $\braket{n}{\hat{\mathcal{O}}}{n}$ and the smooth function $\mathcal{O}(E)$ in the ETH ansatz (1), it is necessary to understand the behavior of $\braket{n}{\hat{\mathcal{O}}_{\perp}}{n}$.
@@
By definition, since $\hat{\mathcal{O}}^{\|}$ is a linear combination of the powers of the Hamiltonian $\hat{H}$, we have $[\hat{H},\hat{\mathcal{O}}^{\|}]=0$. Therefore, the operator $\hat{\mathcal{O}}_{\perp}$ encodes the components of the observable $\hat{\mathcal{O}}$ that do not commute with the Hamiltonian. In particular, when $||[\hat{H},\hat{\mathcal{O}}]||=O(1)$, the norm of $\hat{\mathcal{O}}_{\perp}$ is also of order unity, $||\hat{\mathcal{O}}_{\perp}||=O(1)$. To bound $\braket{n}{\hat{\mathcal{O}}_{\perp}}{n}$ in Eq. (10), we invoke properties of ergodic systems. While $\hat{\mathcal{O}}_{\perp}$ is a nontrivial operator, it bears no special relationship to the eigenvectors |$\ket{n}$ of the Hamiltonian. If the system is ergodic, $\braket{n}{\hat{\mathcal{O}}_{\perp}}{n}$ can be modeled as a diagonal element of a random matrix scaling as $\braket{n}{\hat{\mathcal{O}}_{\perp}}{n}=O(1/\sqrt{D})$ and, therefore, exponentially with the system size, $L$. In this case, Eq. (10) implies that $\braket{n}{\hat{\mathcal{O}}}{n}$ becomes, up to $O(e^{-L})$ corrections, a smooth function of energy, leading to
@@
Figure 5: (a,b) The probability distribution for the site-resolved magnetization, $s_{l}^{z}$, and (c,d) the site-resolved magnetization after removing the approximate microcanonical expectation, $\tilde{s}_{l}^{z}$. Results are shown, respectively, for the HSC model with Random (a,c) and QP (b,d) disorder. For all panels $L=18$, $W=0.5$, with $100$ eigenstates in intervals around $\braket{\epsilon}$ and $\braket{\epsilon}\pm\varepsilon$ for at least $1300$ disorder realizations.
Let us now specialize to the observable of interest, i.e., the site-resolved magnetization, $\hat{s}^{z}_{l}$, and as the Hamiltonian consider $H_{\mathrm{HSC}}$. From Eq. (11), we observe that understanding of the microcanonical expectation value $s^{z}_{l}(E)$ requires evaluation of the coefficients $\gamma_{k}$. Moreover, for a system with near-Gaussian density of states (DoS) and an operator with norm independent of the system size, it was observed [43] that taking only the first $p$ terms in the expansion (9), incurs error decreasing with system size $L$ as $O(1/L^{p-1})$. For this reason, we consider only the first two terms in the expansion $\hat{s}^{z}_{l,\|}=\gamma_{0}\hat{H}_{\mathrm{HSC}}^{0}+\gamma_{1}\hat{H}_{\mathrm{HSC}}^{1}$, resulting in
@@
$\frac{1}{D}\mathrm{Tr}(s^{z}_{l})=\gamma_{0}+\gamma_{1}\frac{1}{D}\mathrm{Tr}(H_{\mathrm{HSC}}),$ (13)
@@
In both steps, we used the fact that $\hat{s}^{z}_{l,\perp}$ is orthogonal to $H^{k}_{\mathrm{HSC}}$ for any integer $k\geq 0$. Directly computing $\mathrm{Tr}(s^{z}_{l})=0$, $\mathrm{Tr}(\hat{H}_{\mathrm{HSC}})/{D}=-1/4$, $\mathrm{Tr}(\hat{H}_{\mathrm{HSC}}\hat{s}_{l}^{z})/{D}=1/4h_{l}-1/(4(L-1))\sum_{l\neq i}h_{i}$ and $\mathrm{Tr}(\hat{H}_{\mathrm{HSC}}^{2})/{D}=(3L^{2}-3L-1)/(16(L-1))+1/4\sum_{i=1}^{L}h_{i}^{2}-1/(4(L-1))\sum_{l\neq j}h_{l}h_{j}$, we use Eq. (13) and (14) to determine the values of $\gamma_{0}$ and $\gamma_{1}$, which, through (11), imply that
@@
Figure 6: The probability distribution of site-resolved magnetization, $s_{l}^{z}$, (a) and the difference of site-resolved magnetization, $\delta s_{l}^{z}$, (b) for disorder-averaged HSC-QP Hamiltonian. The insets for (a) and (b) show the standard deviation with system sizes for $s_{l}^{z}$ and $\delta s_{l}^{z}$, respectively. Panels (c) and (d) show the same curves as in (a) and (b) but in a logarithmic vertical scale. The distributions are for $W=0.5$ considering $100$ eigenstates at the average energy $\braket{\epsilon}$ with at least $4500$ disorder realizations for different system sizes, $L$.
The derived linear relationship between the site-resolved magnetization, $s_{j}^{z}(E)$, and the energy accurately reproduces the behavior of the diagonal matrix element $\braket{n}{\hat{s}^{z}_{l}}{n}$ in the Hamiltonian’s eigenstates. This is illustrated in Fig. 4 which shows that $\braket{n}{\hat{s}^{z}_{l}}{n}$ concentrate, up to fluctuations suppressed exponentially with system size $L$, around the curves $s_{j}^{z}(E)$ at all values of the rescaled energy $\epsilon$. Fitting $\braket{n}{\hat{s}^{z}_{l}}{n}$ as a function of energy, we confirm that the resulting linear fit reproduces the parameters of (15) down to machine precision. In particular, we observe that the zero of $s_{l}^{z}(E)$ occurs at the rescaled energy $\braket{\epsilon}$ corresponding to $\braket{E}$. Further details of the derivations may be found in Appendix A.

### v3

@@ I Introduction
The paper is organized as follows. In Sec. II, we briefly describe the peculiar properties of the probability distribution of the site-resolved magnetization defined as $\braket{n|\hat{s}_{l}^{z}|n}\equiv s_{l}^{z}$, where $l$ is sampled over the site indices, for the Heisenberg spin1/2-1/2- 1 / 2 chain in the presence of the quasiperiodic disorder as encountered in [44], and proceed to explore the distribution of $s_{l}^{z}$ for the Heisenberg spin1/2-1/2- 1 / 2 chain with different types of onsite disorders. This leads us to a conjecture about the relationship between the distribution of the onsite disorder and the properties of the $s_{l}^{z}$ distribution. We show the validity of this conjecture by finding the leading-order approximation for the microcanonical expectation of the site-resolved magnetization in Sec. III, and conclude that the particular structure of the $s_{l}^{z}$ distribution is an artifact of the choice of the energy window, Sec. IV. Finally, in Sec. V, we show that these specific structures have direct consequences for experiments and can be detected in the time dynamics of the $s_{l}^{z}$ expectation values.
Figure 1: The probability distribution of site-resolved magnetization, $s_{l}^{z}$, for disorder-averaged GOE matrices (a) and the HSC-QP Hamiltonian at $W=0.5$ (b) at different system sizes, $L$. For both (a) and (b) panels, 100100100 eigenstates are considered at the rescaled energy $\epsilon=0.5$. The number of disorder realizations is varied, being at least 750075007500 for $L\leq 16$ and 606060 for $L=18$ for panel (a), while for HSC-QP case (b) it is $10^{4}$ disorder realizations for $L\leq 16$ and 250025002500 for $L=18$. The insets show the standard deviations of respective distributions, decreasing exponentially with $L$ for GOE (a) and saturating for HSC-QP model.
Figure 2: The distribution of onsite disorder is depicted for various types of disorders, including random (a), cosine (b), binary (c), tertiary (d), and triangle-wave (e). Additionally, the corresponding distributions of site-resolved magnetization are shown in (f-j), considering 100100100 states near $\epsilon=0.5$ averaged over at least 200200200 disorder realizations for a system of size $L=18$.
@@ II Unusual distribution for the site-resolved magnetization
We begin with a brief description of the distribution of site-resolved magnetization in a Heisenberg spin-1/21/21 / 2 chain (HSC) with quasiperiodic disorder [44]. The HSC is a paradigmatic model extensively employed to study ergodicity breaking and the interplay between disorder and quantum many-body dynamics [45]. The HSC Hamiltonian with open boundary conditions in the presence of an onsite potential reads
@@
where $\hat{s}_{l}^{\alpha}$ are spin-1/21/21 / 2 operators acting on site $l$, with $\alpha\in{x,y,z}$ and $L$ is the system size. Disorder in the model arises from the onsite potential term; initially, we shall consider an example of disorder arising from a quasiperiodic potential (QP[46]. The model has a $U(1)$ symmetry leading to magnetization, $M\equiv\sum_{l=1}^{L}\hat{s}_{l}^{z}$, being conserved. We will restrict ourselves to the magnetization-zero sector in this work. The QP introduces a deterministic, incommensurate modulation in the onsite potential, given by:
@@
where $E$ represents the energy of the state under consideration, while $E_{\min}$ and $E_{\max}$ denote the minimum and maximum eigenvalues in the spectrum, respectively. This definition ensures that $\epsilon$ is normalized to the range [0,1][0,1][ 0 , 1 ].
Figure 3: The distribution of the difference of adjacent onsite disorder, $\delta h_{l}$, is depicted for various types of disorders, including random (a), cosine (b), binary (c), tertiary (d), and triangle-wave (e). Additionally, the corresponding distributions of differences of adjacent site-resolved magnetization, $\delta s_{l}^{z}$, (f-j), considering 100100100 states near $\epsilon=0.5$, averaged over at least 200200200 disorder realizations for a system of size $L=18$.
The observable of interest in this work is the site-resolved magnetization, $\hat{\mathcal{O}}\equiv\hat{s}_{l}^{z}$. In analogy with a random matrix, one could expect the observable to be Gaussian distributed around its mean value and with a standard deviation that is exponentially decreasing with system size. We confirm this expectation numerically by taking the Hamiltonian $\hat{H}$ as a matrix from the Gaussian Orthogonal ensemble [56]. The dimension of $H$ is the same as the Hilbert space dimension of HSC; the diagonal matrix elements of $H$ are taken from the Gaussian distribution $\mathcal{N}(0,1)$ with zero mean and unity variance, while the off-diagonal elements of $H$ are taken from $\mathcal{N}(0,1/{2})$. In Figure 1(a), we show the probability distribution of site-resolved magnetization considering the expectation values $s^{z}_{l}=\braket{n|\hat{s}_{l}^{z}|n}$ in eigenstates |$\ket{n}$ around the rescaled energy $\epsilon=0.5$. We observe a clear single peak structure with the standard deviation decreasing exponentially with system size.
@@
    Tertiary disorder: we consider a distribution of disorder with $h_{l}$ selected randomly, with equal probability, to be $h_{l}\in\{-W,0,W\}$. This model is in the MBL regime for $W_{c}\geq 1.25\sqrt{\frac{3}{2}}$ [62]. In Fig. 2 (d,i) and Fig. 3 (d,i), we show the presence of triple peaks in the $s^{z}_{l}$ distribution, associated with the tertiary distribution of $h_{l}$, while $\delta h_{l}$ has a quinary distribution and correspondingly $\delta s_{l}^{z}$ has 555 peaks with amplitudes decreasing with the distance from $|\delta s_{l}^{z}|=0$.
@@ III Microcanonical expectation for site-resolved magnetization
To proceed more formally, following [43] we define the Hilbert-Schmidt product in operator space as $\braket{\hat{A},\hat{B}}\equiv\frac{1}{{D}}\mathrm{Tr}[\hat{A}^{\dagger}\hat{B}]$, where ${D}$ is the Hilbert space dimension, and the norm as $||\hat{A}||=\sqrt{\braket{\hat{A},\hat{A}}}$. The Hamiltonian of an ergodic many-body system is denoted by $\hat{H}$, and $\hat{\mathcal{O}}$ represents an arbitrary few-body observable. The operator space can be decomposed into two orthogonal subspaces: one spanned by the powers of the Hamiltonian, $\hat{H}^{k}$ (with $k\geq 0$), and an orthogonal complement. Hence, the observable $\hat{\mathcal{O}}$ can be decomposed as
@@
$\braket{n|\hat{\mathcal{O}}|n}=\sum_{k=0}^{\ell}\gamma_{k}E_{n}^{k}+\braket{n|\hat{\mathcal{O}}_{\perp}|n}.$ (10)
To establish a connection between the diagonal matrix element $\braket{n|\hat{\mathcal{O}}|n}$ and the smooth function $\mathcal{O}(E)$ in the ETH ansatz (1), it is necessary to understand the behavior of $\braket{n|\hat{\mathcal{O}}_{\perp}|n}$.
@@
By definition, since $\hat{\mathcal{O}}^{\|}$ is a linear combination of the powers of the Hamiltonian $\hat{H}$, we have $[\hat{H},\hat{\mathcal{O}}^{\|}]=0$. Therefore, the operator $\hat{\mathcal{O}}_{\perp}$ encodes the components of the observable $\hat{\mathcal{O}}$ that do not commute with the Hamiltonian. In particular, when $||[\hat{H},\hat{\mathcal{O}}]||=O(1)$, the norm of $\hat{\mathcal{O}}_{\perp}$ is also of order unity, $||\hat{\mathcal{O}}_{\perp}||=O(1)$. To bound $\braket{n|\hat{\mathcal{O}}_{\perp}|n}$ in Eq. (10), we invoke properties of ergodic systems. While $\hat{\mathcal{O}}_{\perp}$ is a nontrivial operator, it bears no special relationship to the eigenvectors |$\ket{n}$ of the Hamiltonian. If the system is ergodic, $\braket{n|\hat{\mathcal{O}}_{\perp}|n}$ can be modeled as a diagonal element of a random matrix scaling as $\braket{n|\hat{\mathcal{O}}_{\perp}|n}=O(1/\sqrt{D})$ and, therefore, exponentially with the system size, $L$. In this case, Eq. (10) implies that $\braket{n|\hat{\mathcal{O}}|n}$ becomes, up to $O(e^{-L})$ corrections, a smooth function of energy, leading to
@@
Figure 5: (a,b) The probability distribution for the site-resolved magnetization, $s_{l}^{z}$, and (c,d) the site-resolved magnetization after removing the approximate microcanonical expectation, $\tilde{s}_{l}^{z}$. Results are shown, respectively, for the HSC model with Random (a,c) and QP (b,d) disorder. For all panels $L=18$, $W=0.5$, with 100100100 eigenstates in intervals around $\braket{\epsilon}$ and $\braket{\epsilon}\pm\varepsilon$ for at least 130013001300 disorder realizations.
Let us now specialize to the observable of interest, i.e., the site-resolved magnetization, $\hat{s}^{z}_{l}$, and as the Hamiltonian consider $\hat{H}_{\mathrm{HSC}}$. From Eq. (11), we observe that understanding of the microcanonical expectation value $s^{z}_{l}(E)$ requires evaluation of the coefficients $\gamma_{k}$. Moreover, for a system with near-Gaussian density of states (DoS) and an operator with norm independent of the system size, it was observed [43] that taking only the first $p$ terms in the expansion (9), incurs error decreasing with system size $L$ as $O(1/L^{p-1})$. For this reason, we consider only the first two terms in the expansion $\hat{s}^{z}_{l,\|}=\gamma_{0}\hat{H}_{\mathrm{HSC}}^{0}+\gamma_{1}\hat{H}_{\mathrm{HSC}}^{1}$, resulting in
@@
$\frac{1}{D}\mathrm{Tr}(\hat{s}^{z}_{l})=\gamma_{0}+\gamma_{1}\frac{1}{D}\mathrm{Tr}(\hat{H}_{\mathrm{HSC}}),$ (13)
@@
In both steps, we used the fact that $\hat{s}^{z}_{l,\perp}$ is orthogonal to $\hat{H}^{k}_{\mathrm{HSC}}$ for any integer $k\geq 0$. Directly computing $\mathrm{Tr}(s^{z}_{l})=0$, $\mathrm{Tr}(\hat{H}_{\mathrm{HSC}})/{D}=-1/4$, $\mathrm{Tr}(\hat{H}_{\mathrm{HSC}}\hat{s}_{l}^{z})/{D}=1/4h_{l}-1/(4(L-1))\sum_{l\neq i}h_{i}$ and $\mathrm{Tr}(\hat{H}_{\mathrm{HSC}}^{2})/{D}=(3L^{2}-3L-1)/(16(L-1))+1/4\sum_{i=1}^{L}h_{i}^{2}-1/(4(L-1))\sum_{l\neq j}h_{l}h_{j}$, we use Eq. (13) and (14) to determine the values of $\gamma_{0}$ and $\gamma_{1}$, which, through (11), imply that
@@
Figure 6: The probability distribution of site-resolved magnetization, $s_{l}^{z}$, (a) and the difference of site-resolved magnetization, $\delta s_{l}^{z}$, (b) for disorder-averaged HSC-QP Hamiltonian. The insets for (a) and (b) show the standard deviation with system sizes for $s_{l}^{z}$ and $\delta s_{l}^{z}$, respectively. Panels (c) and (d) show the same curves as in (a) and (b) but in a logarithmic vertical scale. The distributions are for $W=0.5$ considering 100100100 eigenstates at the average energy $\braket{\epsilon}$ with at least 450045004500 disorder realizations for different system sizes, $L$.
The derived linear relationship between the site-resolved magnetization, $s_{j}^{z}(E)$, and the energy accurately reproduces the behavior of the diagonal matrix element $\braket{n|\hat{s}^{z}_{l}|n}$ in the Hamiltonian’s eigenstates. This is illustrated in Fig. 4 which shows that $\braket{n|\hat{s}^{z}_{l}|n}$ concentrate, up to fluctuations suppressed exponentially with system size $L$, around the curves $s_{j}^{z}(E)$ at all values of the rescaled energy $\epsilon$. Fitting $\braket{n|\hat{s}^{z}_{l}|n}$ as a function of energy, we confirm that the resulting linear fit reproduces the parameters of (15) down to machine precision. In particular, we observe that the zero of $s_{l}^{z}(E)$ occurs at the rescaled energy $\braket{\epsilon}$ corresponding to $\braket{E}$. Further details of the derivations may be found in Appendix A.